9.18 Given a normal population whose mean is 50 and
whose standard deviation is 5, find the probability
that a random sample of
a. 4 has a mean between 49 and 52.
b. 16 has a mean between 49 and 52.
c. 25 has a mean between 49 and 52.
1
Expert's answer
2020-12-08T08:11:23-0500

"Given \; that, μ=50, σ=5, then,\\
 a) P(49<\bar x<52)=P(\frac{49-50}{\frac{5}{\sqrt{4}}}<Z<\frac{52-50}{\frac{5}{\sqrt{4}}})\\
=P(-0.4<Z<0.8)\\
=P(0<Z<0.4)+P(0<Z<0.8)\\
=0.1554+0.2881=0.4435\\
b) P(49<\bar x<52)=P(\frac{49-50}{\frac{5}{\sqrt{16}}}<Z<\frac{52-50}{\frac{5}{\sqrt{16}}})\\
=P(-0.8<Z<1.6)\\
=P(0<Z<0.8)+P(0<Z<1.6)\\
=0.2881+0.4452=0.7333\\
c) P(49<\bar x<52)=P(\frac{49-50}{\frac{5}{\sqrt{25}}}<Z<\frac{52-50}{\frac{5}{\sqrt{25}}})\\
=P(-1<Z<2)\\
=P(0<Z<1)+P(0<Z<2)\\
=0.3413+0.4772=0.8185\\"